OpenStudy (anonymous):

(sqrt2)/4 plus (sqrt6)/4

OpenStudy (anonymous):

Really, why not just do it? You can't simplify this one much though, it's just \[\frac{\sqrt{2}+\sqrt{6}}{4}\]